Jill traveled against the wind in a small plane for 6 hours. On the return trip traveling with the wind the trip took 4 hours. Find the speed of the wind if the speed of her plane in still air is 90 miles per hour.
Answer by mananth(16949)

Jill traveled against the wind in a small plane for 6 hours. On the return trip traveling with the wind the trip took 4 hours. Find the speed of the wind if the speed of her plane in still air is 90 miles per hour.
..
speed of plane = 90 mph
speed of wind = x mph.
Against wind speed = 90-x
with wind speed = 90+x
..
against wind time = 6 hours
with wind time = 4hours
..
d= rt . d is same to & fro
6(90-x)=4(90+x)
540-6x=360+4x
10x=180
/10
x= 18 mph
...
m.ananth@hotmail.ca